OpenStudy (anonymous):

Simplify: 16 1/2 sqrt 100

OpenStudy (anonymous):

you mean \[(16 1/2) \sqrt{100}\] ?

OpenStudy (anonymous):

\[16.5\sqrt{100}\]

OpenStudy (anonymous):

\[16.5\sqrt{100}\] = 16.5 * 10 = 165

OpenStudy (anonymous):

im sorry, i wrote it wrong. can u explain to mean how to simplify 16 1/2?

OpenStudy (anonymous):

you can not further simplify sixteen and a half - it is in its simplest form you could write it as a decimal though, is that what you want? 16 1/2 as a decimal is 16.5

OpenStudy (anonymous):

thank you so much, now I c why I was confused.

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

now I c why I couldnt get the lowest common factor

OpenStudy (anonymous):

its simplified at its lowest form

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

any more questions?

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

np

OpenStudy (anonymous):

yes, how about simplifying sqrt 100? square roots confuses me

OpenStudy (anonymous):

it is simple, a square root of 100 is simply a number that after being multiplied by itself gives 100, in this case 10

OpenStudy (anonymous):

explain it to me, so that I could do my next problem on my own

OpenStudy (anonymous):

but i want to see how you came up with your answer. thank u

OpenStudy (anonymous):

ok wait a second...

OpenStudy (anonymous):

wanna become my fan?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

:)